# Soft-CAM: Making black box models self-explainable for high-stakes decisions

### v1

@@ 4 Experimental setup
We evaluated our approach on three publicly available medical datasets spanning three imaging modalities: the Kaggle Diabetic Retinopathy (DR) [20], Retinal OCT [31], and the RSNA Chest X-Ray (CXR) [1]. The first dataset comprised high-resolution retinal color fundus images, each labeled with a DR severity score ranging from 00 (No DR) to $4$ (Proliferative DR). The second dataset included retinal OCT B-scans images categorized into Drusen, Diabetic macular edema, Choroidal neovascularization, and Normal cases. The final dataset consisted of high-resolution frontal-view chest radiographs labeled for pneumonia detection, with bounding box annotations for pneumonia cases. Additionally, lesion annotations were obtained for $65$ DR images from the Kaggle dataset [17] and $40$ Drusen images from the retinal OCT dataset [16]. Each dataset was split into training, validation, and test sets using different dataset-specific train-validation-test proportions, ensuring that all samples from the same patient were assigned to the same split to prevent data leakage. For full details, see Appendix A.1.
@@
The effectiveness of our method was evaluated using two widely used black-box CNN architectures: ResNet-50 [25] and VGG-16 [49]. They differ primarily in the design of their classification heads, where ResNet employs a single fully connected layer, while VGG uses multiple. In both models, we explicitly replaced the classification head with our convolutional evidence map layer, adapting the architecture to enable interpretability (see Appendix A.2). The models were sourced from Torchvision [36], initialized with pre-trained weights from ImageNet, and fine-tuned using a consistent setup [^1] Our code with datasets is available at https://anonymous.4open.science/r/SoftCAM-E1A3/. Training was performed over $70$ epochs with a mini-batch size of $16$ on an NVIDIA A40 GPU using PyTorch [42]. A range of data augmentation and preprocessing techniques was applied (see Appendix A.3). For complete training details, see Appendix A.4.
Baseline CAM-based methods.
We qualitatively and quantitatively assessed the explanations generated by our method (SoftCAM) against post-hoc explanation techniques from several state-of-the-art class attribution map-based methods, applied to their respective black-box models. Specifically, we compared our approach with gradient-based methods, including GradCAM [47], Integrated Gradient (Itgd Grad.) [53], Guided Backpropagation (Guided BP) [51], as well as gradient-free methods such as ScoreCAM [56] and LayerCAM ([30]). Guided BP and Itgd Grad. have consistently performed well in producing saliency maps for explaining black-box CNN classifiers on retinal images [8, 17], while GradCAM has shown strong localization performance for chest X-ray interpretation [46]. Gradient-based methods were implemented from Captum [34], whereas gradient-free methods were implemented via TorchCAM [21]. For full descriptions of these methods, see Appendix A.5.
@@
Models were evaluated on both classification performance and explainability. For binary tasks, performance was measured using accuracy and AUC, while for multi-class tasks, accuracy and the quadratic Cohen’s kappa score were used. AUC reflects class separability, whereas the kappa score captures agreement beyond chance. To assess explainability, we employed several quantitative metrics: Top-k localization precision [18], activation precision [9, 43], activation consistency [18], and faithfulness [16, 39]. We further extended activation precision to define activation sensitivity. For full descriptions of the explainability metrics, see Appendix B.
@@ 5 Results
5.1 Making black box CNNs self-explainable maintains classification performance
Table 1: Classification performance for binary disease detection on the test sets. We denote the SoftCAM versions of ResNet and VGG with a .
@@ 5.1 Making black box CNNs self-explainable maintains classification performance
Binary Multi-class Binary Multi-class Binary
Acc. AUC Acc. $\kappa$ Acc. AUC Acc. $\kappa$ Acc. AUC
VGG-16 $0.907$ $0.938$ $0.863$ $0.835$ $0.994$ $1.0$ $0.967$ $0.955$ $0.952$ $0.989$
dense VGG $0.915$ $0.942$ $0.861$ $0.834$ $0.994$ $1.0$ $0.963$ $0.947$ $0.957$ $0.999$
sparse VGG $0.911$ $0.938$ $0.859$ $0.827$ $0.988$ $0.999$ $0.947$ $0.929$ $0.953$ $0.990$
ResNet-50 $0.899$ $0.923$ $0.850$ $0.800$ $0.994$ $0.999$ $0.970$ $0.963$ $0.953$ $0.988$
dense ResNet $0.899$ $0.926$ $0.851$ $0.811$ $0.994$ $1.0$ $0.974$ $0.960$ $0.942$ $0.983$
sparse ResNet $0.895$ $0.923$ $0.851$ $0.801$ $0.996$ $1.0$ $0.963$ $0.955$ $0.941$ $0.979$
We first evaluated our method on clinically relevant classification tasks, including retinal disease classification from color fundus and OCT retinal images, as well as pneumonia detection from chest X-rays. For the fundus and OCT retinal datasets, both binary classification ({0} vs. {1-4} for fundus and Normal vs. Drusen for OCT) and multi-class classification tasks were considered, as reference labels were available. In contrast, the RSNA CXR dataset only included labels for pneumonia detection, restricting the evaluation to the binary task. For each CNN architecture, the “dense” model corresponds to our method without regularization ($\lambda_{1}=\lambda_{2}=0$), while the “sparse” model is obtained by applying a lasso penalty ($\lambda_{2}=0$) and choosing an appropriate value for $\lambda_{1}$ (e.g. $\lambda_{1}=1.10^{-6}$ for VGG and $\lambda_{1}=5.10^{-5}$ for ResNet on the fundus dataset). The sparsity parameter was selected by balancing classification accuracy and AUC on the validation set (see Appendix C).
Our results show that SoftCAM models, which use explicit self-explainable class evidence maps, preserve classification performance comparable to their corresponding black-box counterparts (Tab. 1). Moreover, introducing the Lasso regularization penalty on the class evidence map did not significantly degrade performance; in some cases, it even led to slight improvement. These findings suggest that using convolutional layers in the classification head is an effective and promising approach for developing high-performing, self-explainable CNN models.
@@ 5.2 SoftCAM provides inherently interpretable visual explanations
Figure 2: Example explanations generated by different methods from ResNet-50. The first column shows disease images with reference annotations, indicated by green markers or bounding boxes. Each row, from top to bottom, corresponds to fundus, OCT, and Chest X-ray images, respectively. The next five columns present saliency maps generated by post-hoc explanation methods, gradient-free (ScoreCAM, LayerCAM) and gradient-based (GradCAM, Guided BP, Itgd Grad). The final two columns showcase our proposed inherently interpretable dense and sparse SoftCAM explanations.
We qualitatively compared the evidence maps of SoftCAM variants with saliency maps generated by the five state-of-the-art CAM-based methods. Overall, our method produced more visually interpretable maps with high evidence regions centered on annotated lesions (Fig. 2). We observed that the regions highlighted by the sparse SoftCAM models are mostly a subset of those identified by the dense SoftCAM, reflecting the effect of the sparsity constraint. Additional results, including those for VGG-16 and other illustrative examples, are provided in Appendix D.1.
On healthy images, sparse SoftCAM evidence maps exhibited overall more negative activations, in contrast to the positive activations observed on disease images. To assess this quantitatively, we computed the activation consistency [18], calculating the proportion of positive and negative activations across disease and healthy samples. These findings were consistent with the qualitative observations (e.g. dense vs. sparse SoftCAM on the fundus dataset using ResNet: $0.55$ vs. $0.27$ for the proportion of positive activation on disease images). For full analysis, see Appendix D.2.
@@ 5.3 SoftCAM provides localized and faithful explanations
Figure 3: Quantitative evaluation of explanations generated by different methods. The first row shows the localization precision of the saliency maps on the Fundus and OCT datasets, evaluated against their respective ground truth. The second row presents the sensitivity analysis assessing the faithfulness of the generated explanations. Columns a,b show ResNet results, while c,d correspond to VGG. Higher precision means better localization; lower sensitivity implies more reliable explanations.

### v2

@@ 3 Experimental setup
We evaluated our approach on three publicly available medical datasets spanning three imaging modalities: the Kaggle Diabetic Retinopathy (DR) [kaggle_dr_detection], Retinal OCT [kermany2018identifying], and the RSNA Chest X-Ray (CXR) [rsna_dataset]. The first dataset comprised high-resolution retinal color fundus images labeled with DR severity score ranging from 0 (No DR) to $4$ (Proliferative DR). The second dataset included retinal OCT B-scans images labeled for three retinal disease conditions. The final dataset consisted of high-resolution frontal-view chest radiographs labeled for pneumonia detection, with bounding boxes for pneumonia cases. Additionally, clinicians annotated lesions in $65$ DR images from the Kaggle dataset [djoumessi2024inherently] and $40$ drusen images from the retinal OCT dataset [djoumessi2024actually] to allow directly clinical evaluation. Each dataset was split into training, validation, and test sets, ensuring that all samples from a given patient remained in the same split. For full details, see Appendix B.
@@
Our method was evaluated using two widely used black-box CNN architectures: ResNet-50 [he2016deep] and VGG-16 [simonyan2014very]. They differ primarily in the design of their classification heads, where ResNet employs a single fully connected layer, while VGG uses multiple layers. More details on the training setup2, including data preprocessing and augmentation, are provided in the Appendix C.
Post-hoc baseline.
SoftCAM was compared with widely used post-hoc methods (Appendix D), including CAM gradient-based approaches such as GradCAM [selvaraju2017grad] and LayerCAM [jiang2021layercam]; the CAM gradient-free method ScoreCAM [wang2020score]; and backpropagation-based techniques such as Integrated Gradients (Itgd Grad) [sundararajan2017axiomatic] and Guided Backpropagation (Guided BP) [springenberg2014striving]. Each method was evaluated on its respective black-box CNN models.
@@
The models were evaluated for predictive performance and explainability. For explainability, we used several quantitative metrics, including top-k localization precision [donteu2023sparse]; activation precision [barnett2021case], faithfulness [yeh2019fidelity], and activation consistency [donteu2023sparse], and we further extended activation precision to define activation sensitivity. Together, these metrics quantify both alignment with expert clinician annotations and alignment with the model’s actual decision-making process. Full metric descriptions are provided in Appendix E.
@@ 4 Results
4.1 Making black box CNNs explainable maintains classification performance
Table 1: Classification performance for binary disease detection on the test sets. SoftCAM variants of both CNNs are denoted by SC, with $\ell_{\lambda}$ indicating the applied penalty.
@@ 4.1 Making black box CNNs explainable maintains classification performance
Acc. AUC Acc. AUC Acc. AUC
VGG-16 $0.907$ $0.938$ 0.994 1.0 $0.952$ $0.989$
VGG-16SC $\mathbf{0.915}$ 0.942 0.994 1.0 0.957 0.999
$\ell_{1}$-VGG-16SC $0.911$ $0.938$ $0.988$ $0.999$ $0.953$ $0.990$
ResNet-50 0.899 $0.923$ $0.994$ $0.999$ 0.953 0.988
ResNet-50SC 0.899 0.926 $0.994$ 1.0 $0.942$ $0.983$
$\ell_{1}$-ResNet-50SC $0.895$ $0.923$ 0.996 1.0 $0.941$ $0.979$
We first evaluated our method on clinically relevant binary classification tasks, including retinal disease classification from color fundus ({0} vs. {1-4}) and OCT retinal images (Normal vs. Drusen), as well as pneumonia detection from chest X-rays, using accuracy and AUC as primary metrics. For each CNN architecture, the $\text{}^{\text}{\ltxml@oqmark@open\textquotedblleft\penalty 10000\thinspace\thinspace SC\textquotedblright\ltxml@oqmark@close{}}$ model corresponds to our method without regularization ($\lambda_{1}=\lambda_{2}=0$), while its $\ell_{1}$ variant is obtained by applying a lasso penalty ($\lambda_{2}=0$) and choosing an appropriate value for $\lambda_{1}$ (e.g. $\lambda_{1}=1.10^{-6}$ for VGG and $\lambda_{1}=5.10^{-5}$ for ResNet on the fundus dataset). The sparsity parameter was selected by balancing classification accuracy and AUC on the validation set (Appendix F).
Our results show that SoftCAM-based models (Tab. 1), with explicit class evidence maps, preserve classification performance comparable to their corresponding black-box baselines. Moreover, introducing the lasso regularization on the class evidence map did not significantly degrade performance; in some cases, it even led to a slight improvement. These findings suggest that using convolutional layers in the classification head is an effective and promising approach for developing high-performing, self-explainable CNN models.
@@ 4.2 SoftCAM provides inherently interpretable visual explanations
We qualitatively compared the evidence maps of the SoftCAM variants with the saliency maps generated by the five state-of-the-art CAM and back-propagation-based methods. Overall, our method produced more visually interpretable maps with high evidence regions centered on annotated lesions (Fig. 2). We observed that the regions highlighted by the sparse SoftCAM models are typically a subset of those identified by the base SoftCAM, reflecting the effect of the sparsity constraint in reducing irrelevant activations. Additional results, including those for VGG-16 and other examples, are provided in Appendix G.1.
On healthy images, the sparse SoftCAM evidence maps exhibited overall more negative activations, in contrast to the positive activations observed on disease images. To assess this quantitatively, we computed the activation consistency score [donteu2023sparse], calculating the proportion of positive and negative activations in disease and healthy samples from the test set. These findings were consistent with qualitative visualization (e.g. sparse SoftCAM vs. SoftCAM on the fundus dataset using ResNet: $0.55$ vs. $0.27$ for the average proportion of positive activations on disease images). For a full analysis, see Appendix G.2.
Figure 2: Example explanations generated by different methods from ResNet-50. The first column shows disease images with reference annotations (green markers or bounding boxes). Row from top to bottom corresponds to fundus, OCT, and Chest X-ray images, respectively. The next five columns present saliency maps from different post-hoc explanation methods. The last two columns show our proposed inherently interpretable dense and sparse SoftCAM explanations.
@@ 4.3 SoftCAM provides localized and faithful explanations
Figure 3: Quantitative evaluation of explanations methods. The first row shows the localization precision of saliency maps on fundus and OCT datasets. The second row presents the sensitivity analysis assessing faithfulness. Columns a,b show ResNet results, and c,d correspond to VGG. Higher precision indicates better localization; lower sensitivity reflects more reliable explanations.
To quantitatively assess the explanations provided by our SoftCAM-based evidence maps compared to post-hoc saliency methods, we first evaluated their localization precision, which measures how consistently the highlighted regions in the explanation maps align with clinician-annotated disease findings. Following [donteu2023sparse], we computed the Top-k (k=15) localization precision by upsampling each explanation map to the input resolution, splitting it into non-overlapping $33\times 33$ patches, and calculating the proportion of positively activated patches that overlap with ground truth annotations. Although inherently interpretable, SoftCAM-based explanations performed competitively overall in terms of localization precision (Fig. 3). Notably, the sparse SoftCAM with the ResNet backbone outperformed all other methods with the highest top-k precision (Appendix G.3 and G.4), and ranked second only in top-3 precision on the fundus dataset (Fig. 3a), behind Guided BP, which benefits from high-resolution saliency maps. Furthermore, the base and sparse SoftCAM typically achieved higher precision with fewer top-k regions, leading to fast convergence, particularly on the Fundus and OCT datasets. This suggests that sparse SoftCAM more consistently highlights fewer, yet truly relevant regions, whereas post-hoc methods produce broader and less specific activations, resulting in higher false-positive rates.
Subsequently, we evaluated the faithfulness (also referred to as sensitivity) of the evidence maps generated by our SoftCAM-based explanations in comparison to post-hoc saliency maps. Sensitivity analysis evaluates how much the highly activated regions in an explanation map contribute to the model’s prediction [yeh2019fidelity], thereby assessing whether the highlighted areas actually influence the model’s decision-making process. To do this, we split the input images into non-overlapping $33\times 33$ patches, then progressively removed the top-ranked patches (based on attribution scores) and measured the relative change in model confidence. We conducted this evaluation on samples that were correctly predicted by both the black-box CNNs and their corresponding SoftCAM variants from the test sets. We found that the sparse SoftCAM generally outperformed other methods, notably on the OCT and RSNA datasets (Fig. 3; Appendix G.3 and G.4). On the fundus dataset, both the base and sparse SoftCAM models performed slightly below the best-performing post-hoc methods, with Guided BP yielding the highest sensitivity scores, followed by Integrated Gradients (Fig. 3). On the OCT dataset, the SoftCAM variants outperformed all post-hoc methods with the ResNet model, while with the VGG the sparse and base versions ranked second and third, respectively. On the RSNA dataset, the sparse SoftCAM achieved the highest sensitivity, surpassing all other methods, with the base SoftCAM ranking second with ResNet and third with VGG (Appendix G.4).
Since the CXR dataset provided larger bounding boxes localizing disease regions, unlike the point-wise lesion annotations available in the fundus and OCT datasets, we computed activation precision [barnett2021case], which measures the proportion of the class-guided explanation that falls within the ground-truth bounding boxes, emphasizing precision by penalizing only false positives. However, it does not account for sensitivity by not penalizing false negatives. To address this, we extended this metric to activation sensitivity (Appendix E.3), which penalizes false negatives to better assess the explanation completeness, especially important in clinical imaging tasks where missing relevant regions can be critical. We further investigated how different regularization methods qualitatively and quantitatively affect explanations. While lasso regularization promoted sparsity by shrinking most false positive activations to zeros, ridge regularization encouraged small (but nonzero) values, resulting in denser evidence maps. To evaluate this, we trained a ridge SoftCAM model ($\lambda_{1}=0$) and compared its performance to the sparse SoftCAM, as well as to the post-hoc explanation methods. The ridge regularization strength was selected to balance predictive performance ($\lambda_{2}=7.10^{-5}$ vs. $\lambda_{2}=2.10^{-4}$ for ResNet and VGG; Appendix H.1).
[…]
4.5 SoftCAM provides faithful explanations for multi-class tasks
We extended our method to multi-class settings, training the SoftCAM models from ResNet and VGG for DR detection (5 classes, Kaggle dataset) and retinal disease classification (4 classes, OCT dataset). The training setup remained consistent with the binary task, adjusting only the number of classes in the evidence layer and selecting appropriate lasso penalties. Detailed classification results, qualitative visualizations, and faithfulness analyses of the multi-class SoftCAM-based variants and their baseline are provided in the Appendix I. Briefly, both unregularized and sparse models achieved test accuracies comparable to their baseline (Acc $\approx 0.85$ vs. $0.97$ on Fundus and OCT), with SoftCAM, particularly the sparse variant, providing the most interpretable class-wise explanations with the best sensitivity.
5 Discussion
Here, we introduced SoftCAM, a lightweight architectural modification that makes standard black-box CNNs inherently interpretable without relying on post-hoc explanations. SoftCAM replaces the final pooling and FCL with 1×1 convolutions, producing class-specific evidence maps that directly inform predictions. This design also supports ElastiNet regularization on the evidence maps to improve explanations. As a result, SoftCAM produces aligned predictions and explanations in a single forward-pass, yielding resource-efficient, self-explainable CNNs. SoftCAM-based variants were validated on ResNet-50 and VGG-16 across three medical imaging modalities—fundus photographs, retinal OCT scans, and Chest X-rays—achieving classification performance comparable to their black-box baselines. For explainability, SoftCAM was benchmarked against five state-of-the-art post-hoc saliency methods—ScoreCAM, LayerCAM, GradCAM, Guided BP, and Itgd Grad—using qualitative and quantitative metrics that capture both alignment with human knowledge and alignment with the model’s true decision-making, demonstrating superior interpretability without compromising predictive performance. In particular, localization precision and activation sensitivity assess how well explanations match clinically relevant biomarkers using expert annotations, while faithfulness measures how highlighted regions truly influence the model’s predictions. Interestingly, our results revealed a discrepancy between human-aligned and model-aligned metrics (Fig. 3): explanations with the best faithfulness did not always yield the strongest alignment with expert annotations. This suggests that models may rely on clinically hidden signals to humans and highlights the need for new evaluation metrics that jointly account for human knowledge and models’ internal reasoning.
Despite its promising results, SoftCAM has some limitations that warrant further investigation. First, SoftCAM relies on the final low-resolution feature maps (e.g., 16×16 for standard CNNs), limiting the spatial granularity needed for fine-grained medical tasks that require lesion- or pixel-level precision. In our case, both ResNet and VGG models use large receptive fields, producing low-resolution feature maps. Consequently, the class-evidence layer generates coarse-grained explanations, though ElasticNet constraints help improve localization. Future work could focus on generating higher-resolution explanations with softCAM to overcome coarseness from downsampling in deep CNNs; integrating SoftCAM with other architectures like ViTs [dosovitskiy2020image]; extending it to tasks such as weakly supervised segmentations or object detection; and evaluating its utility in other high-stake domains, including agriculture or medical modalities such as dermoscopy for skin cancer detection or MRI for brain tumor detection. Finally, SoftCAM was compared only against a subset of relevant post-hoc methods, consistent with our primary goal of avoiding post-hoc explanations for CNNs classifiers. Future work could extend this evaluation to include additional post-hoc techniques and self-explainable models, such as part-prototype networks [chen2019looks], concept-based models [koh2020concept], or attention-based explainable architectures [djoumessi2025hybrid].
[…]
This project was supported by the Hertie Foundation, the German Science Foundation (Excellence Cluster EXC 2064 “Machine Learning—New Perspectives for Science”, project number 390727645; BE 5601/14-1, project number 571331899). The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Kerol Djoumessi.
Appendix A Related work
Explainable AI (XAI) methods for image analysis can be categorized into attribution and non-attribution-based approaches [he2022survey, hossain2023explainable]. Attribution-based methods aim to explain “where” important features are located by generating saliency maps that assign importance scores to pixels or regions. In contrast, non-attribution methods focus on explaining “why” a decision was made, typically providing global explanations using approaches such as prototype- or concept-based explanations [chen2019looks, koh2020concept]. Attribution-based methods usually provide post-hoc local explanations for black-box models after training. In contrast, non-attribution-based methods are generally inherently interpretable by design [chen2019looks, brendel2018bagnets, koh2020concept, bohle2022b], embedding transparency within their architecture. However, inherently interpretable models are typically model-agnostic and often face reduced performance, reflecting a persistent trade-off between interpretability and predictive performance [rudin2019stop]. To bridge this gap, we propose SoftCAM, a protocol that makes traditional black-box CNNs inherently interpretable. Our approach builds on and generalizes prior work.
aubreville2019transferability proposed a dual-branch architecture in which the first branch performed classification using a black-box CNN, while the second branch generated post-hoc explanations, requiring two separate forward passes. In the explanation branch, the global average pooling (GAP) layer is removed and the linear classifier is replaced with convolutional layers that share weights with the main classifier during inference, producing class-specific activation maps. In contrast, SoftCAM integrates both prediction and explanation within a single, end-to-end forward pass, eliminating the need for redundant computation and weights sharing. Similarly, donteu2023sparse leveraged a convolutional classifier to generate explicit class-evidence maps and applied a sparsity constraint to enhance interpretability in a self-explainable Bag-of-Local-Features model (BagNet) brendel2018bagnets. SoftCAM extends and generalizes this concept to standard black-box CNNs, broadening its applicability to various medical imaging tasks and modalities.
Appendix B Datasets
We evaluated our approach on three publicly available medical imaging datasets spanning three different modalities: the Kaggle Diabetic Retinopathy (DR) kaggle_dr_detection, the Retinal OCT dataset kermany2018identifying, and the RSNA Chest X-ray (CXR) dataset rsna_dataset.
  • Kaggle DR Dataset. This dataset comprises $88,702$ high-resolution retinal fundus images labeled for DR severity on a 5-point scale from 0 (No DR) to $4$ (Proliferative DR). After applying an automated quality filtering pipeline using an ensemble of EfficientNet models trained on the ISBI2020111https://isbi.deepdr.org/challenge2.html challenge dataset, we retained $45,923$ images from $28,984$ subjects. The resulting class distribution was $73\%$, $15\%$, $8\%$, $3\%$, and $1\%$ for classes 0-4 respectively. For binary classification (early DR detection), we grouped class {0} vs. {1,2,3,4}, yielding an imbalance of $73\%$ vs. $27\%$. Additionally, lesion annotations for $65$ images were obtained from djoumessi2024inherently for evaluating the model’s explanations at localizing DR-related lesions.
  • Retinal OCT Dataset. This dataset consists of $108,315$ B-scans categorized into four classes: Drusen, Diabetic macular edema (DME), Choroidal neovascularization (CNV), and Normal. A separate test set of $1,000$ B-scans is provided. Following djoumessi2024actually, we excluded low-resolution scans (width $\leq$ 496). As preliminary experiments showed that using the full dataset did not significantly improve performance, we subsampled the training set (by randomly removing half of the healthy images following djoumessi2024actually) to $34,962$ scans ($8,616$ Drusen, $26,346$ Normal) for binary classification (Drusen vs. Normal), preserving the original class imbalance ($73\%$ vs $27\%$). Additionally, we used $40$ drusen-annotated B-scans from djoumessi2024actually to evaluate the model’s explanations at localizing drusen lesions. For the multi-class classification task, the training was randomly reduced to $17,200$ images while maintaining the original class distribution: $45\%$ Normal, $34\%$ CNV, $10\%$ DME, and $9\%$ Drusen.
[…]
Appendix C Implementation details
C.1 Baseline models
The effectiveness of our method was evaluated using two widely adopted black-box CNN architectures: ResNet-50 and VGG-16. These models were chosen due to their distinct architectures, such as depth, theoretical receptive field size, and classification head design, which allow for a broad assessment of our method’s generalizability. In both models, the standard classification head was systematically replaced with our proposed convolution-based evidence map layer to enable inherent interpretability.
For ResNet50, we removed the global average pooling layer and final fully connected layer (FCL), substituting them with a class evidence layer consisting of $C$ convolutional filters ($1\times 1$, stride 1), where $C$ is the number of output classes. This layer directly produces class-specific evidence maps (Sec. 2.2). For VGG-16, whose classifier head consist of several fully connected layers, each FCL was replaced by an equivalent $1\times 1$ convolutional layer. Specifically, an FCL of size $b_{1}\times b_{2}$ was transformed into a convolutional layer of size $b_{1}\times b_{2}\times 1\times 1$, preserving the original parameter count and model capacity. These architectural changes maintain model complexity and capacity while introducing interpretability directly into the classification mechanism.
C.2 Data preprocessing and augmentation
Fundus images were preprocessed by cropping them to a square shape using a circle-fitting method as described in Mueller_fundus_circle_cropping. All datasets were then resized to $512\times 512$ pixels, except for the retinal OCT dataset, which was resized to $496\times 496$ to better match its original lower resolution. Image intensities were normalized using the mean and standard deviation computed from the respective training sets.
During training, standard transformations were applied across all datasets. These included flipping, rotation, random cropping, and translation, each applied with a fixed probability. For the Kaggle dataset, which contains color fundus images, additional color augmentations were introduced to improve generalization.
C.3 Training setup
[…]
Appendix D Baseline post-hoc methods
SoftCAM was benchmarked against widely used post-hoc methods, spanning CAM-based and backpropagation-based approaches. CAM-based methods produce coarse-grained explanations due to their reliance on low-resolution convolutional feature maps, whereas backpropagation-based techniques provide fine-grained pixel-level attributions that perverse full spatial resolutions. CAM-based methods include gradient-based approaches such as GradCAM and LayerCAM, as well as the gradient-free method ScoreCAM. Backpropagation-based techniques include Integrated Gradients and Guided Backpropagation. Gradient-based methods primarily differ in how they aggregate gradients to compute importance weights, while gradient-free methods vary in how these weights are estimated without backpropagation.
Guided Backpropagation and Integrated Gradients have consistently performed well in generating saliency maps to explain black-box CNN classifiers on retinal images ayhan2022clinical; djoumessi2024inherently, while GradCAM has shown strong localization performance for chest X-ray interpretation saporta2022benchmarking. Below is a brief description of the five post-hoc methods used.
CAM
ScoreCAM
wang2020score. A gradient-free method that eliminates the need for gradient information by assessing the importance of each activation map based on its forward-pass contribution to the target class score, and produces the final output via a weighted sum of these maps.
LayerCAM
jiang2021layercam. A gradient-based method that generates class activation maps by leveraging the element-wise product of ReLU-activated gradients and feature maps at any convolutional layer, enabling fine-grained, spatially precise visual explanations without requiring global average pooling.
GradCAM
selvaraju2017grad. A gradient-based approach that uses the gradients of the target class flowing into the final convolutional layer to produce a coarse localization map, highlighting important regions in the image by upsampling the resulting map.
Guided backpropagation (Guided BP)
springenberg2014striving. A gradient-based approach that modifies the standard backpropagation process to propagate only positive gradients through positive activations, producing fine-grained visualizations that highlight features strongly activating specific neurons in relation to the target output.
Integrated Gradient (Itgt Grad.)
sundararajan2017axiomatic. A gradient-based method that attributes model predictions to input features by computing the path integral of gradients along a straight-line path from a baseline to the actual input, yielding fine-grained explanations.
ScoreCAM and LayerCAM were implemented with TorchCAM torcham2020, while the other methods were implemented from Captum kokhlikyan2020captum.
Appendix E Explainability metrics
For binary tasks, performance was evaluated using accuracy and AUC, while for multi-class tasks, accuracy and the quadratic Cohen’s kappa score ($\kappa$) were used. The AUC measures class separability, whereas the kappa score captures the agreement beyond chance.
Explainability was assessed using several quantitative metrics, including activation consistency donteu2023sparse, top-k localization precision donteu2023sparse, activation precision barnett2021case, further extended to activation sensitivity, and faithfulness yeh2019fidelity. Together, these metrics quantify both alignment with expert clinical knowledge (top-k localization, activation precision, and activation sensitivity) and alignment with the model’s true decision-making process (faithfulness).
E.1 Activation consistency
The activation consistency donteu2023sparse measures how well local explanations (i.e., positive or negative activation in attribution maps) reflect the true disease or healthy labels across a dataset. Intuitively, an interpretable model should consistently show positive activation regions associated with pathology for disease samples, while producing minimal or negative activations for healthy samples. This metric therefore assesses whether explanation maps globally align with the semantic meaning implied by ground-truth labels.
Following donteu2023sparse, activation consistency is computed as the proportion of positive activations in the attribution maps of disease samples and the proportion of negative activations in the attribution maps of healthy samples, averaged over the test set. A higher score indicates more coherent and label-aligned explanations.